### v1

# Relations Between the Inequality Indices Gini, Pietra and Kolkata: Theory and Data Analysis

Abstract
We study here relations between three inequality indices, namely the Gini ($g$), Pietra ($p$) and Kolkata ($k$) introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function $L(x)$ introduced in 1905. The Kolkata index (which corresponds to a fixed point of the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$) gives the fraction of wealth $k$ possessed by the richest 1k1-k fraction of people ($k$ = 0.8 corresponds to Pareto’s 80-20 law from 1896). We show rigorously that the Pietra or Robin Hood index $p$ should equal to the excess wealth fraction 2k12k-1 possessed by the richest 1k1-k fraction of people. Our numerical data analysis for US IRS Income data (1983-2022), Bollywood (India) movie income data (1999-2024) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always greater than unity but the deviation is never more than five percent. Assuming some simple analytic form for the Lorenz function, we also derived the relations $k=(1/2)+(3/8)g$ for small $g$ values and $p/g=3/4$. However, these relations generally deviate significantly for larger $g$ or $k$ values when compared with observations.
I Introduction
Quantitative measurements of economic inequalities, specifically the income or wealth inequalities, in different societies or countries started with the introduction of Lorenz function Lorenz1905 in 1905. The Lorenz function or curve is represented graphically by plotting the cumulative fraction of the income or wealth ($L(x)$) earned or possessed by the $x$ fraction of the poorest individuals, when the population of the society is arranged in the ascending order of their income (from now on income will mean wealth of any kind of agent in countries, societies, institutions etc.; see Fig. 1). When everyone earns the same, the Lorenz curve becomes the dotted-dashed (black) diagonal in Fig. 1. Because of the inequalities among the agents (and their ordering from the poorest to the richest), the typical shape of the Lorenz curve will have the form of the red line (with $L(0)=0$ and $L(1)=1$). Although $L(x)$ contains all the information of inequality, typical values of the inequality indices help summarizing or characterizing the nature of inequality in the society. One of the oldest and still most popular one is the Gini index Gini1912 ($g$) given by the normalized area between the equality line and the Lorenz curve (see Fig. 1). The Pietra index Pietra15 ($p$), also known as Allen22 Robin Hood index or Hoover index or Schulz index, is given by the maximum vertical distance between the equality line and the Lorenz curve (see Fig. 1). This index (length value) corresponds to the percentage of the total population’s income that would have to be redistributed to make everyone’s income equal.
Figure 1: The Lorenz curve or function ($L(x)$, in red) shows the proportion of total wealth owned by a fraction ($x$) of people in ascending order of wealth. The black dotted-dashed line represents a scenario of perfect equality in which everyone possesses the same amount of wealth. The Gini index ($g$) is calculated from the area between the Lorenz curve and the equality line (shaded region), normalized by the total area (= 1/2) under the equality line. The Pietra index $p$ given by the length of the maximum vertical distance of the Lorenz curve from the equality line and indicated here by the solid vertical line in black. The complementary Lorenz function ($L_{c}(x)\equiv 1-L(x)$ is) shown in green. The Kolkata index ($k$) is determined by the fixed point of the complementary Lorenz curve: $L_{c}(k)=k$ or L(k)=1kL(k)=1-k. Geometrically it gives the point at which the Lorenz curve intersects the diagonal line perpendicular to the equality line and it gives the fraction $k$ of wealth that is possessed by the richest 1k1-k fraction of people. As such $k=0.8$ corresponds to Pareto’s 80-20 law.
In 2005 Hirsch noted Hirsch05 that the success in citations of the papers by the scientists could be quantified by an index $h$ (known as Hirsch index today) corresponding to the fixed point of the nonlinear (Ziff-like) citation function ($f(c)$) giving the number $f$ of papers having citation $c$ : $h=f(h)$. Soon this index became a very popular one to quantify the success inequalities among the scientists. Inspired by that, and also noting that the nonlinear Lorenz function $L(x)$ has trivial fixed points at $x=0$ and $x=1$, we defined Ghosh14 the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$, which has a nontrivial fixed point at $k$ such that $L_{c}(k)=k$, where $k\geq 1/2$ (see Fig. 1). This $k$ is called the Kolkata index (see also Banerjee20 ; Banerjee23 ) to epitomize the extreme economic inequality in this Indian city. A simple geometric argument for the Lorenz function then suggests that 1k1-k fraction of the rich people possess $k$ fraction of the total wealth. As such it generalizes then the Pareto’s 80-20 law Pareto1896 where $k=0.80$. Soon it was observed (see e.g., Banerjee23 ; Manna22 ) that in physical systems with self-organized competitive dynamics among the constituent degrees of freedom the inequality levels (in all cluster or avalanche size distributions) tend to grow towards a universal level $g=k\simeq 0.87$ (somewhat above the Pareto value $k=0.80$) and remains thereafter in the self-organized critical state of the system. Later, assuming Joseph22 a minimal polynomial expansion of the Lorenz function $L(x)$, we could derive a simple linear relationship between the Gini and Kolkata indices growing up to $g=k=0.80$. This linear relationship of Kolkata ($k$) with Gini ($g$) indices agreed generally with our data analysis (see Banerjee23B ) for lower values of $g$.
Figure 2: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for US economy income data (IRS data irs ; Yakovenko22 for the period 1982-2022). The values of $k$ and $g$ are growing with time, because of increasing rate of withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate a value higher than unity as predicted by a rigid theoretical argument (relation (1)). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from the theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table I of the Appendix.
Figure 3: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for US economy income tax return data (IRS data irs ; Yakovenko22 for the period 1982-2022). The values of $k$ in the tax data (which can be argued to represent the prevailing inequality status better) seems to have grown a little beyond the Pareto value ($k=0.80$) with increasing withdrawal of public welfare programs. The upper left and lower right insets, showing the values of p/(2k1)p/(2k-1) against $g$ and $k$ respectively, indicate values slightly higher than unity (as predicted by a rigid theoretical argument; see relation (1). The lower right inset shows the values of p/g against years and show values considerably different from 3/4, as obtained from the theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the values of the inequality indices and of their relations are given in Table II of the Appendix.
II Theoretical analysis of Inequality indices
We will first explore analytically if there are any relations among these three indices, namely Gini ($g$), Pietra ($p$) and Kolkata ($k$). To begin with, we note that since by construction the Kolkata index ($k$) is given by the fraction of the total wealth possessed by the 1k1-k fraction of richest people ($k>1/2$; $k=1/2$ corresponds to equality), they possess precisely an extra amount of 2k12k-1 fraction of total wealth. The Pietra index Allen22 (or Robin Hood index) $p$ then should equal to this excess amount, suggesting
p=2k1.p=2k-1. (1)
In order to find theoretically the other plausible relations among those indices, let us proceed with a (Landau-like; see e.g., Manna22 ) minimal polynomial expansion of the Lorenz function:
$L(x)=Ax+Bx^{2},$ (2)
where $A>0$ and $B>0$, such that $L(x)$ becomes a monotonically increasing function of $x$ and $A+B=1$ to ensure $L(x)=1$ for $x=1$ (also ensuring $L(x)=0$ at $x=0$). One can then express the Gini index through
$g=1-2\int_{0}^{1}L(x)dx=1-A-(2/3)B=B/3.$
The Pietra index $p$ is given by the maximum value of the difference between $x$ (equality line) and $L(x)$ (Lorenz curve) for variation of $x$ within 0 to 1:
$p=max[x-L(x)]=max[B(x-x^{2}]=B/4,$
giving,
$p=(3/4)g.$ (3)
Also noting that the Kolkata index $k$ is given by the fixed point equation $L_{c}(k)=k$ or the complementary Lorenz function $L_{c}(x)\equiv 1-L(x)$, giving (see also Manna22 ; Banerjee23 )
$k=1/2+(3/8)g,$ (4)
for $g\rightarrow 0$.
Figure 4: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for movie income inequalities (Bolloywood India data 13bbollywood for the period 1999-2024). The values of $k$ in the tax data (which can be argued to represent the prevailing inequality status better) seems to have grown a little beyond the Pareto value ($k$ = 0.80) with increasing withdrawal of public welfare programs. The upper left inset, showing the values of p/(2k1)p/(2k-1) against the years, indicate values higher than unity as predicted by a rigid theoretical argument (relation (1)). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the values of the inequality indices and and of their relations are given in Table III of the Appendix.
III Numerical Data Analysis
We will now analyze the extensive data sets from various income data sources (US IRS for the forty year period 1983 to 2022 irs ; Yakovenko22 , Bolywood Box Office income data for the period 1999 to 2024 13bbollywood ), comparing how the above relations (2), (3) and (4) fit the results of our data analysis. We will analyze extensive citation data (from google scholar google-scholar ) of the prolific Nobel laureate research scientists (physicists, chemists, medicinal scientists and economists) for the years 2020 to 2025.
Figure 5: Growths of Pietra ($p$) and Kolkata ($k$) indices against Gini ($g$) index values for the citation inequalities among the papers published by different Nobel Laureates in economics, physics, chemistry and medicine during the last six years (Google Scholar open-access data google-scholar for scientists, each having their own home page with verified e-mail address and more than 100 papers). The values of $k$ (and also of $g$ for the inequalities in citation distributions across the publications of individual Nobel laureates) have grown, because of extreme competition among the scientists, quite beyond the Pareto value ($k$ = 0.80, and closer to self-organized critical system value Manna22 ; Banerjee23B ) because of any public welfare kind of support system (for the producers) and the market being quite competitive. The upper left and lower right insets, showing the values of p/(2k1)p/(2k-1) against $g$ and $k$ respectively, indicate values slightly higher than unity (as predicted by a rigid theoretical argument; see relation (1). The lower right inset shows the values of $p/g$ against years and show values considerably different from 3/4, as obtained from the theoretical relation (3) (obtained an additional assumption (2) of the minimal form of the Lorenz function). Details of the inequality indices and of their relations are given in Table IV of the Appendix.
IV Summary and Conclusion
As we discussed, all the three inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) are derived from the Lorenz function $L(x)$ (see Fig. 1). $2g$ is given by the area between the equality line and the Lorenz curve. $p$ is given by the maximum value of xL(x)x-L(x) and is interpreted as a measure of the fraction of money the ‘rich’ people should transfer to the ‘poor’ people (neither rich or poor are specifically defined here), so that a redistribution would bring equality (Robin Hood index Allen22 ). $k$ is given by the fixed point of the Complementary Lorenz Function $L_{c}(x)\equiv 1-L(x)$: $L_{c}(k)=k$. Geometrically (see Fig. 1) it gives the fraction $k$ of wealth (citations) possessed by clearly defined 1k1-k fraction of rich people (papers). We therefore argued the excess wealth of the rich is precisely equal to 2k12k-1, which if they transfer to the poor $k$ fraction of people, then (on redistribution) equality would be achieved. As such p=2k1p=2k-1 (relation (1)).
As we see from numerical data analysis (see Appendices I-IV) and upper-left insets of Figs. 2 - 5, the $p$ is a little higher (at most $5\%$ off) from that suggested by relation (1). Next we assumed a minimal analytic and polynomial form (2) for the Lorenz function $L(x)$ and calculated $g$ and $p$ from there, giving $p=(3/4)g$ (relation (3)). As we see from the tables I-IV and Figs. 2-5, there are large deviations from relation (2). Finally, solving for $k$ from (2) and (3), in the small $g$ limit, we got $k=(1/2)+(3/8)g$ (relation (3)). Numerical data analysis for the validity of this relation are indicated in each of the main parts and bottom right insets of Figs. 2-5.
We studied here relations between three inequality indices, namely the Gini ($g$) Gini1912 , Pietra ($p$) Pietra15 and Kolkata ($k$) Ghosh14 ; Banerjee20 ; Banerjee23 introduced in 1912, 1915 and 2014 respectively and all are derived from the Lorenz function Lorenz1905 $L(x)$ introduced in 1905. We showed rigorously that the Pietra or Robin Hood index $p$ should equal to the excess wealth 2k12k-1 possessed by the richest 1k1-k people, giving relation (1). Our numerical data analysis results for US IRS Income data (1983-2022; see Tables I, II of the Appendix, and Figs. 2,3), Bollywood (India) movie income data (1999-2024; see Table III of the Appendix and Fig. 4) and the citation inequalities across the publications by forty Nobel Laureates (2020-2025; see Table IV of the Appendix and Fig. 5) in Economics, Physics, Chemistry and Medicine clearly shows that p/(2k1)p/(2k-1) is always slightly greater than unity and the deviation is never more than five percent. Assuming some simple analytic form (2) for the Lorenz function (cf. Joseph22 ), we also derived here the relations (3) $p/g=3/4$ and (4) $k=(1/2)+(3/8)g$ for small values of $g$. However, these relations generally deviate significantly for larger $g$ and $k$ values (see Tables I-IV and Figs. 2-5) when compared with observations.
Acknowledgement
We are grateful to Suchismita Banerjee and Shohini Sen for their valuable comments and suggestions.
Appendix: Detailed numerical estimates of inequality indices from data analysis
Table 1: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) values for the income inequalities from movie for U.S. IRS Income Data (1983–2022) datasets collected from ref. irs ; Yakovenko22 . For each year, the total income collections of individuals are analyzed.
Year Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1983 0.47418 0.34280 0.66894 1.0146 0.7229 0.7126 1984 0.47711 0.34289 0.67076 1.0040 0.7187 0.7158 1985 0.48136 0.34672 0.67253 1.0048 0.7203 0.7168 1986 0.49412 0.35602 0.67640 1.0091 0.7205 0.7140 1987 0.51365 0.36991 0.68403 1.0050 0.7202 0.7166 1988 0.53394 0.38312 0.69106 1.0026 0.7175 0.7157 1989 0.53094 0.38250 0.68970 1.0082 0.7204 0.7146 1990 0.52898 0.38110 0.68854 1.0107 0.7204 0.7128 1991 0.52472 0.37906 0.68734 1.0117 0.7224 0.7141 1992 0.53185 0.38256 0.68960 1.0089 0.7193 0.7130 1993 0.53100 0.38191 0.68960 1.0071 0.7192 0.7141 1994 0.53198 0.38212 0.69071 1.0018 0.7183 0.7170 1995 0.54026 0.38946 0.69434 1.0020 0.7209 0.7194 1996 0.55188 0.39894 0.69943 1.0002 0.7229 0.7227 1997 0.55998 0.40463 0.70153 1.0039 0.7226 0.7198 1998 0.56631 0.40807 0.70376 1.0013 0.7206 0.7196 1999 0.57604 0.41659 0.70780 1.0024 0.7232 0.7215 2000 0.58418 0.42321 0.71140 1.0010 0.7245 0.7237 2001 0.56116 0.40616 0.70285 1.0011 0.7238 0.7230 2002 0.55009 0.39857 0.69902 1.0013 0.7246 0.7236 2003 0.55508 0.40325 0.70161 1.0001 0.7265 0.7264 2004 0.57073 0.41385 0.70533 1.0078 0.7251 0.7195 2005 0.58713 0.42413 0.71052 1.0073 0.7224 0.7171 2006 0.59445 0.42721 0.71290 1.0033 0.7187 0.7163 2007 0.60078 0.43191 0.71543 1.0024 0.7189 0.7172 2008 0.58395 0.42027 0.70950 1.0030 0.7197 0.7175 2009 0.56321 0.40960 0.70280 1.0099 0.7273 0.7202 2010 0.57349 0.41558 0.70664 1.0056 0.7247 0.7206 2011 0.57799 0.41882 0.70862 1.0038 0.7246 0.7219 2012 0.59759 0.43488 0.71601 1.0066 0.7277 0.7229 2013 0.58345 0.42418 0.71113 1.0045 0.7270 0.7237 2014 0.59374 0.43475 0.71569 1.0078 0.7322 0.7265 2015 0.59442 0.43629 0.71649 1.0076 0.7340 0.7284 2016 0.58799 0.43158 0.71424 1.0072 0.7340 0.7287 2017 0.59553 0.43759 0.71800 1.0036 0.7348 0.7321 2018 0.59473 0.43703 0.71829 1.0010 0.7348 0.7341 2019 0.59072 0.43407 0.71669 1.0016 0.7348 0.7336 2020 0.60100 0.44124 0.72022 1.0018 0.7342 0.7328 2021 0.62792 0.46403 0.73005 1.0085 0.7390 0.7327 2022 0.60345 0.44444 0.72064 1.0072 0.7365 0.7313
Table 2: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) values for the tax-return inequalities from U.S. IRS Income Tax Data (1983–2022) datasets collected from ref. irs ; Yakovenko22 . For each year, the total tax collections of individuals are analyzed.
Year Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1983 0.66335 0.48721 0.74086 1.0114 0.7345 0.7262 1984 0.66753 0.48835 0.74363 1.0022 0.7316 0.7299 1985 0.67246 0.49438 0.74581 1.0056 0.7352 0.7311 1986 0.69308 0.51074 0.75262 1.0109 0.7369 0.7290 1987 0.70073 0.51685 0.75674 1.0066 0.7376 0.7328 1988 0.71296 0.52519 0.76215 1.0017 0.7366 0.7354 1989 0.70412 0.51864 0.75899 1.0013 0.7366 0.7356 1990 0.70247 0.51917 0.75930 1.0011 0.7391 0.7383 1991 0.70487 0.52181 0.76089 1.0001 0.7403 0.7402 1992 0.71938 0.53487 0.76601 1.0054 0.7435 0.7396 1993 0.72833 0.54241 0.76992 1.0048 0.7447 0.7412 1994 0.72995 0.54421 0.77123 1.0032 0.7455 0.7431 1995 0.73822 0.55293 0.77611 1.0013 0.7490 0.7480 1996 0.74976 0.56455 0.78198 1.0010 0.7530 0.7522 1997 0.75437 0.56893 0.78243 1.0072 0.7542 0.7488 1998 0.76372 0.57613 0.78651 1.0054 0.7544 0.7503 1999 0.77301 0.58454 0.79122 1.0036 0.7562 0.7535 2000 0.77856 0.59549 0.79470 1.0103 0.7649 0.7570 2001 0.76491 0.57785 0.78725 1.0058 0.7554 0.7511 2002 0.77313 0.58737 0.79221 1.0050 0.7597 0.7559 2003 0.77373 0.58709 0.79204 1.0052 0.7588 0.7549 2004 0.78684 0.60559 0.80048 1.0077 0.7696 0.7638 2005 0.79995 0.62076 0.80998 1.0013 0.7760 0.7750 2006 0.80278 0.62305 0.81109 1.0014 0.7761 0.7750 2007 0.80582 0.62481 0.81229 1.0004 0.7754 0.7751 2008 0.80036 0.62046 0.80949 1.0024 0.7752 0.7734 2009 0.80642 0.63018 0.81371 1.0044 0.7815 0.7780 2010 0.81101 0.63382 0.81576 1.0036 0.7815 0.7787 2011 0.79568 0.61789 0.80758 1.0044 0.7766 0.7731 2012 0.80744 0.62826 0.81359 1.0017 0.7781 0.7768 2013 0.80558 0.62664 0.81260 1.0023 0.7779 0.7761 2014 0.81165 0.63613 0.81674 1.0042 0.7837 0.7805 2015 0.80929 0.63505 0.81648 1.0033 0.7847 0.7821 2016 0.80136 0.62672 0.81236 1.0032 0.7821 0.7796 2017 0.80401 0.63025 0.81495 1.0006 0.7839 0.7834 2018 0.81284 0.64080 0.81722 1.0100 0.7883 0.7805 2019 0.80787 0.63630 0.81471 1.0109 0.7876 0.7791 2020 0.82292 0.65008 0.82091 1.0129 0.7900 0.7799 2021 0.82025 0.64352 0.81709 1.0147 0.7845 0.7732 2022 0.81013 0.62775 0.81029 1.0116 0.7749 0.7660
Table 3: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) values for the inequalities from movie income datasets containing year-wise box office earnings (1999-2022) for Bollywood (India) movies, collected from ref. 13bbollywood . For each year, the total collections of individual films are recorded. Only years with more than 10 released movies are included to ensure reliable statistical representation. Here NN represent the number of produced and released movies in each year.
Year Tot. No. of movies Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g 1999 52 0.41093 0.29326 0.64469 1.0134 0.7136 0.7042 2000 51 0.41637 0.29554 0.64453 1.0224 0.7098 0.6942 2001 51 0.55679 0.41486 0.70681 1.0030 0.7451 0.7429 2002 51 0.36364 0.26156 0.63000 1.0060 0.7193 0.7150 2003 52 0.45222 0.33932 0.66963 1.0002 0.7503 0.7502 2004 51 0.44658 0.34524 0.67206 1.0033 0.7731 0.7706 2005 52 0.38679 0.29895 0.64882 1.0044 0.7729 0.7695 2006 50 0.52425 0.41003 0.70370 1.0065 0.7821 0.7771 2007 51 0.49243 0.36600 0.68287 1.0007 0.7433 0.7427 2008 51 0.49477 0.36377 0.68170 1.0010 0.7352 0.7345 2009 64 0.61159 0.46545 0.72633 1.0283 0.7610 0.7401 2010 140 0.77307 0.63066 0.80724 1.0263 0.8158 0.7949 2011 123 0.78239 0.65696 0.81366 1.0472 0.8397 0.8018 2012 133 0.78140 0.64864 0.81088 1.0432 0.8301 0.7957 2013 137 0.76403 0.60528 0.79360 1.0308 0.7922 0.7686 2014 146 0.80382 0.65775 0.81760 1.0355 0.8183 0.7902 2015 165 0.79784 0.65836 0.81791 1.0355 0.8252 0.7969 2016 216 0.82970 0.68791 0.83025 1.0415 0.8291 0.7961 2017 252 0.85300 0.70442 0.84432 1.0229 0.8258 0.8073 2018 219 0.84407 0.71848 0.84777 1.0330 0.8512 0.8240 2019 244 0.84798 0.72439 0.85467 1.0212 0.8543 0.8365 2020 76 0.86551 0.73448 0.85860 1.0241 0.8486 0.8286 2021 50 0.79775 0.63359 0.81054 1.0201 0.7942 0.7785 2022 165 0.84379 0.69608 0.84151 1.0191 0.8249 0.8095 2023 227 0.89157 0.77006 0.87826 1.0179 0.8637 0.8485 2024 230 0.89728 0.76409 0.87516 1.0184 0.8516 0.8362
Table 4: Gini-index ($g$), Pietra-index ($p$) and Kolkata-index ($k$) for the citation inequalities among the papers published by each of the Nobel Laureates (collected, for years 2020-2025, from Google Scholar google-scholar ). Here $h$ represents the Hirsch index Hirsch05 for the individual scientists listed.
Categories Year Name of Nobel laureates Tot. no. of papers Tot. no. of citations Hirsch index ($h$) Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g Economics 2020 Paul Milgrom 398 118799 83 0.9123 0.7965 0.8941 1.0105 0.8731 0.8640 2020 Robert Wilson 259 34825 58 0.8799 0.7379 0.8685 1.0012 0.8386 0.8376 2021 David Card 695 103999 117 0.8957 0.7627 0.8785 1.0075 0.8515 0.8452 2021 Guido Imbens 393 119042 102 0.8869 0.7522 0.8751 1.0025 0.8481 0.8459 2021 Josh Angrist 436 107740 92 0.9212 0.7891 0.8922 1.0060 0.8566 0.8515 2022 Ben Bernanke 717 147600 118 0.9210 0.7984 0.8984 1.0021 0.8669 0.8651 2022 Philip H. Dybvig 139 55939 41 0.9426 0.8339 0.9139 1.0073 0.8846 0.8782 2024 Daron Acemoglu 1358 277596 183 0.9167 0.7928 0.8939 1.0064 0.8649 0.8594 2024 James Robinson 845 131696 102 0.9465 0.8293 0.9146 1.0001 0.8761 0.8760 2024 Simon Johnson 947 95164 64 0.9693 0.8962 0.9477 1.0008 0.9246 0.9238 2025 Joel Mokyr 471 32139 71 0.9049 0.7548 0.8747 1.0071 0.8341 0.8282 2025 Philippe Aghion 647 146570 145 0.8733 0.7205 0.8586 1.0046 0.8250 0.8212 Physics 2020 Roger Penrose 639 101446 102 0.9022 0.7727 0.8859 1.0011 0.8565 0.8555 2021 Giorgio Parisi 1127 111103 136 0.8437 0.6694 0.8346 1.0002 0.7934 0.7933 2021 Syukuro Manabe 298 49325 89 0.8291 0.6687 0.8282 1.0186 0.8065 0.7918 2022 John F. Clauser 138 22474 33 0.9450 0.8298 0.9137 1.0029 0.8781 0.8756 2022 Aspect Alain 772 41611 76 0.9342 0.7984 0.8972 1.0050 0.8546 0.8503 2023 Anne LHuillier 514 36201 86 0.8470 0.6792 0.8375 1.0063 0.8018 0.7968 2023 Ferenc Krausz 1122 90527 130 0.8860 0.7335 0.8644 1.0064 0.8279 0.8226 2024 Geoffrey Hinton 751 975672 192 0.9440 0.8310 0.9155 1.0001 0.8803 0.8802 2024 John Hopfield 306 94961 95 0.8964 0.7381 0.8687 1.0009 0.8234 0.8227 2025 John Clarke 1081 50125 111 0.8501 0.6913 0.8409 1.0140 0.8132 0.8019 2025 Michel Devoret 754 78163 133 0.8586 0.7110 0.8478 1.0220 0.8281 0.8102 Chemistry 2020 Emmanuelle Charpentier 269 64602 65 0.9309 0.8030 0.9014 1.0002 0.8626 0.8624 2020 Jennifer Doudna 864 156101 164 0.8582 0.6868 0.8429 1.0014 0.8003 0.7992 2021 Benjamin List 337 49436 103 0.7482 0.5753 0.7868 1.0029 0.7690 0.7667 2021 David MacMillan 578 87442 134 0.8288 0.6704 0.8296 1.0171 0.8089 0.7952 2022 Carolyn Bertozzi 1021 103476 153 0.8081 0.6306 0.8132 1.0067 0.7804 0.7752 2022 Morten Meldal 412 32092 68 0.8261 0.6382 0.8165 1.0083 0.7726 0.7663 2023 Louis Brus 468 92751 122 0.8175 0.6606 0.8274 1.0088 0.8081 0.8010 2023 Moungi G. Bawendi 989 178538 189 0.8323 0.6666 0.8314 1.0057 0.8009 0.7964 2024 Demis Hassabis 171 238294 100 0.8530 0.6991 0.8492 1.0009 0.8196 0.8189 2024 David Baker 2690 201046 227 0.8431 0.6868 0.8399 1.0102 0.8146 0.8064 2025 Omar M. Yaghi 741 275500 197 0.8478 0.6895 0.8427 1.0061 0.8133 0.8084 2025 Susumu Kitagawa 1129 104460 146 0.7994 0.6226 0.8112 1.0002 0.7788 0.7787 Medicine 2020 Michael Houghton 546 61896 107 0.8429 0.6749 0.8366 1.0025 0.8007 0.7987 2022 Svante Paabo 587 151556 179 0.7805 0.6006 0.7989 1.0046 0.7696 0.7660 2023 Katalin Karikó 226 32576 66 0.8429 0.6874 0.8408 1.0084 0.8155 0.8087 2024 Victor Ambros 179 73309 72 0.8827 0.7315 0.8653 1.0011 0.8287 0.8278 2025 Shimon Sakaguchi 676 143255 136 0.8794 0.7323 0.8638 1.0065 0.8327 0.8274